3-DIMENSIONAL POTENTIAL ENERGY SURFACE OF THE Ar-SH COMPLEX

\underline{YOSHIHIRO SUMIYOSHI} and YASUKI ENDO, Department of Basic Science, Graduate School of Arts and Sciences, The University of Tokyo, Komaba, Meguro-ku, Tokyo 153-8902, Japan.

All the pure rotational transitions reported in the previous studies,^{a}^{,}^{b} and recently observed rotation vibration transitions, P = 1/2 - 3/2, for Ar-SH^{c} and Ar-SD have been simultaneously analyzed to determine an intermolecular potential energy surface of Ar-SH in the ground state. In the analysis, the SH vibration was explicitly considered to calculate ro-vibrational energies of the complex on a 3-dimensional intermolecular potential energy surface (3D-IPS), where ab initio calculations at the RCCSD(T)/aug-cc-pVTZ + BF level of theory have been performed to facilitate the results as initial values for the least-squares analysis. The determined 3D-IPS was able to fit all the pure-rotational and ro-vibrational data for Ar-SH and Ar-SD without introducing empirical correction terms required in the previous study.^{b}

aY. Sumiyoshi, et al. J. Chem. Phys. \underline{113}, 10121 2000.

bY. Sumiyoshi, et al. J. Mol. Spectrosc. \underline{222}, 22 2003.

cY. Sumiyoshi, et al. Talk TH11, 59th Ohio State University International Symposium on Molecular Spectroscopy, 2004.